OpenStudy (anonymous):

inverse cosine of square root of 3 over 2? cos-1(square root of 3/2)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

cos(pi/6)=square of 3/2

OpenStudy (anonymous):

\[\text{ArcCos}\left[\frac{\sqrt{3}}{2}\right]=\frac{\pi }{6}\] Am not sure the square root of what was.

myininaya (myininaya):

what was?

myininaya (myininaya):

oops i missed my word root above

OpenStudy (anonymous):

was the argument of the cosine (square root 3)/2 or square root (3/2)

myininaya (myininaya):

i assumed it to be over just the numerator

OpenStudy (anonymous):

Thanks for responding.

myininaya (myininaya):

lol everyone needs love

myininaya (myininaya):

rob look at the post before this? do you agree with my attachment?

myininaya (myininaya):

the post by galaxy

myininaya (myininaya):

or i mean thread

OpenStudy (anonymous):

not sure what thread you are referring to. do you have the web address?

myininaya (myininaya):

ok you see all the questions right? right underneath this thread by hanbru there is galaxy (if sinx=1/3 and secy=5/4, where x and y lie btw 0 and pi/2, blah blah